OpenStudy (anonymous):

dividing powers a^3b^2 c/a^3 b c^5

OpenStudy (anonymous):

help i will give a medal

OpenStudy (anonymous):

Hey! So, what we're talking about here are exponent laws. The law we can see here is a version of the division law.. \[\frac{ x^m }{ x^n }=x ^{m-n}\] (Note: this only works if the numerator and the denominator of the fraction only have multiplication AND when the base of the exponents is the same.) For example.. \[\frac{ x^5 }{ x^3 }\] We can just subtract the two exponents to get..\[x^2\]If the values of the exponents were switched (if the three was on top and the five was on the bottom) then the reciprocal would be true..\[\frac{ 1 }{ x^2 }\]In your case we have to do this a few times. \[\frac{ a^3b^2c }{ a^3bc^5 }\]Both of your a's have a power of 3 so when we subtract they become 0. We only lose one of the b's. Finally, the exponent of the c on the bottom loses one because of the c on top leaving us with..\[\frac{ b }{ c^4 }\]Hope I helped! P.S. if you need more info on the exponent laws you can probably find it here: http://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (anonymous):

thank you